a litre of orange fruit drink contains 22%orange juice.how many millilitres of orange juice must be added to produce a mixture containing 50% orange juice?
Answer by mananth(16949)

--------percent ---------------- quantity
Orange drink 22 ---------------- 1000 ml
Orange juice 100 ---------------- x ml
Mixture 50 ---------------- 1000 + x ml

22*1000 +100x =50(1000+x)
22000+100x=50000+50x
100x+-50x=50000-(-22000)
50 x = 28000
/ 50
x = 560
560 ml of Orange juice